x represents number of girls among 10 children. Use range rule of thumb to determine whether 1 girl in 10 births is significantly low. Give maximum and minimum


1
Expert's answer
2021-05-24T02:17:02-0400

The range rule of thumb tells that the range is generally about four times the standard

deviation.

"\mu = E(X) = \sum xP(x)"

"E(X) = 0\times 0.003+1 \times 0.016+2 \times 0.036+--------+10 \times 0.024 = 5.104"

Standard Deviation = 1.769

The Range Rule of Thumb says that the range is about four times the standard deviation.

"R = 4 \sigma"

Maximum value "= \mu + 2\sigma = 5.104+2 \times 1.769 = 8.642"

Minimum value "= \mu-2 \sigma = 5.10-2 \times 1.769 = 1.566"